An investment of $1,000 will return $60 annually forever. What is its internal rate of return? Please show work.

A.

Cannot be determined.

B.

6%

C.

0.60%

D.

16.67%

E.

60%
Answer #1

IRR is the rate of return at which NPV is zero.

NPV = Value of Investment - PV of cash flows expected in future (which is a perpetuity in this case)

PV of a perpetuity is

PV of Perpetuity = – PV = Present Value D = Dividend or Coupon per period r = discount rate

- 0001 = 0

1000 = 60/r

r = 60/1000 = 6%